OpenStudy (anonymous):

Find f(6) if f(x) = x 2(to the second power) ÷ 3 + x. 4 10 18 18 is the answer idk why I even posted this srry

OpenStudy (nurali):

f(x)=x^2/3+x f(6)=(6)^2/3+6 f(6)=36/9 f(6)=4

OpenStudy (anonymous):

Thanks! @Nurali

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

ok @Nurali